AnimeGhoul8863:

Let f(x) = 2x − 6. Solve f−1(x) when x = 2.

AnimeGhoul8863:

\[f(x) = 2x - 6\] \[f(x)=2(x) - 6\] \[y=2(x)-6\] \[x=2y-6\] \[+6 +6\] \[\frac{ 2-6 }{ 2 }\] \[\frac{ 2y }{ 2 }\] \[\frac{ 2-4 }{ 2 } = f^1(x)\]

AnimeGhoul8863:

^how did i do so far

AnimeGhoul8863:

@TheSmartOne

TheSmartOne:

Made a mistake You added 6 but then when you went to the next line you're subtracting it

AnimeGhoul8863:

wait huh let me look

TheSmartOne:

And then after that line I'm not sure what happened

AnimeGhoul8863:

i see i was looking at my notes and put what the note said

AnimeGhoul8863:

\[\frac{ 2+6 }{ 2 }\] \[\frac{ 2y }{ 2 }\] \[\frac{ 8 }{ 2 } = f^-1(x) \]

AnimeGhoul8863:

^fixed?

AnimeGhoul8863:

what do i have to fix now

TheSmartOne:

I don't get what 2y/2 has to do with anything but the rest is correct

AnimeGhoul8863:

its suppose to be 2+6 over 2 = 2y over 2

TheSmartOne:

Oh that makes sense